Whereas the choice of a comparison date has no effect on answer obtained with compound interests, the same cannot be said of simple interest. Find the amount to be paid at the end of 11 years which is equivalent to two payments of $200 each, the first paid immediatly and the second to be paid at the end of 6 years. Assume 6% simple interest is earned from the date each payment is made and use a comparison date of

a) The end of 11 years. ANSWER = $ .

b) The end of 15 years. ANSWER = $ .
Answer #1

Future value = Present value * (1 + RT)

a)

x/(1 + 0.06*11) = 200 + 200/(1+0.06 *6)

amount to paid at end of 11 years x = 576.11

b)

x/(1 + 0.06*15) = 200 + 200/(1+0.06 *6)

amount to paid at end of 15 years x = 659.41